### v2

# Towards Identifiability of Interventional Stochastic Differential Equations

Abstract
We study identifiability of stochastic differential equation (SDE) models under multiple interventions. Our results give the first provable bounds for unique recovery of SDE parameters given samples from their stationary distributions. We give tight bounds on the number of necessary interventions for linear SDEs, and upper bounds for nonlinear SDEs in the small noise regime. We experimentally validate the recovery of true parameters in synthetic data, and motivated by our theoretical results, demonstrate the advantage of parameterizations with learnable activation functions.
1 Introduction
High-throughput perturbations, e.g., using CRISPR systems, coupled with single-cell genomic sequencing provide the opportunity to study the dynamic and causal relationships (known as the “gene regulatory network”, GRN) between genes at an unprecedented resolution, under a wide set of possible interventions, informing which genes should be targeted for treatment [12]. While substantial progress has been made in applying machine learning and causal inference methods to such genomic data, there has been limited theoretical work to understand when such models are identifiable.
Due to the high degree of gene co-expression, substantial noise and unmeasured modalities (e.g., chromatin state, protein expression), latent confounding is particularly severe in genomic data, leading to considerable false positive rates among popular methods [20]. Causal disentanglement aims to learn causal factors in spite of these confounders, mainly focusing on directed acyclic graph (DAG) based methods. To demonstrate these methods are well-founded, there is considerable effort devoted to understanding which models have identifiability guarantees [23]. However, these models suffer from inherent weakness, in particular 1) being unable to represent cycles (a common occurrence in GRNs) or 2) approximate continuous-time dynamical models.
There has been renewed interest in inferring GRNs directly with dynamical systems [31]. There is precedent for this type of modeling to represent the so- called “Waddington landscape” [41], the hypothetical energy surface that cells traverse as they differentiate. Furthermore, SDEs are commonly used for simulating transciptomic datasets from a given gene regulatory network [32, 11], to benchmark GRN recovery algorithms.
As recently expounded in the context of diffusion models (which can be thought of as time-discretized SDEs), SDEs can fit observational data [39]. But without a restriction on the model, it is impossible to identify the true SDE that will generalize to new interventions, either on unseen gene perturbations or intervening on several genes simultaneously. Furthermore, dynamical systems typically learn from many trajectories. But the destructive nature of sequencing makes it impossible to observe the trajectory of a single cell at multiple time-points, and in general it is difficult to obtain any time-series genomics data due to the high expense. Therefore, practitioners often only collect data at the end of an experiment, i.e., from the stationary distribution of the system. Consequently, several dynamics-based models rely on single time points for evaluation, and this lack of trajectories may render the true biological model unidentifiable.
Our interest in this work is to verify which parametric assumptions are necessary for dynamical systems to have these guarantees. Namely:
How many interventions are necessary to identify the parameters of a stochastic differential equation, only given access to the stationary distribution?
Contributions
In this work, we offer the first analysis of identifiability of stochastic differential equations with interventions. Specifically:
  • We characterize tight bounds on the number of interventions necessary for identifiability of linear SDEs with shift interventions.
  • We extend this analysis to nonlinear SDEs in the small noise regime, showing that identifiability is possible even without knowing the activation function of the true model.
  • We apply this insight to synthetic and semi-synthetic data to confirm the efficacy of learned activations in causal SDEs, which improve expressiveness without sacrificing a simple structure from which to extract a gene regulatory network.
2 Setup
2.1 Notation
We will write the elementary basis vectors as $\{e_{i}\}_{i}$. We will consider $\sigma:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ as any elementwise function, i.e. $\sigma_{i}(x)=\sigma_{i}(x_{i})$. This includes elementwise activations as they are applied in multilayer perceptrons (MLPs), but we also allow for elementwise functions where each component acts differently. Writing $\sigma^{\prime}$ will, unless otherwise described, denote the map $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ that applies elementwise the derivative of each component function, i.e. $\sigma^{\prime}(x)=[\sigma_{1}^{\prime}(x_{1}),\dots,\sigma_{n}^{\prime}(x_{n})]$. We will use $Jf$ to denote the Jacobian of a vector valued function $f$, and $\Delta f$ to denote the Laplacian of $f$.
[…]
2.2 Stochastic Differential Equations
We consider SDEs of the following form, where $B_{t}$ is standard Brownian motion:
$dX_{t}=v(X_{t})dt+\sqrt{\epsilon}dB_{t}.$ (1)
We will only consider autonomous systems, i.e. where the drift and noise terms have no dependence on time $t$. We will enforce the weak conditions on the drift and noise to guarantee a unique stationary distribution [4], with a density $p$ that satisfies the Fokker-Planck equation,
$0=-\nabla\cdot(pv)+\frac{\epsilon}{2}\Delta p.$ (2)
2.3 Linear SDEs
We need some classical facts about linear SDEs, which are better understood than their nonlinear cousins, since Fokker-Planck can be solved explicitly.
Theorem 2.1 ([36]).
Consider the SDE
$dX_{t}=(LX_{t}+c)dt+QdB_{t}$ (3)
Assume $L$ is Hurwitz, i.e., all its eigenvalues have strictly negative real parts, and $Q$ is full rank. Then the unique stationary distribution is $\mathcal{N}(-L^{-1}c,\omega)$ where $\omega$ is the unique solution to the Lyapunov equation,
$L\omega+\omega L^{T}+QQ^{T}=0.$ (4)
The covariance may also be written as a matrix integral:
[…]
2.4 Interventional SDEs
We focus on the setting where we only observe the SDE through the induced stationary density under $\mathcal{I}$ different shift interventions. Specifically, there are vectors $\{c_{i}\}_{i=1}^{\mathcal{I}}$ with each $c_{i}\in\mathbb{R}^{n}$, and we observe the stationary distribution of the SDE,
$dX_{t}=(v(X_{t})+c_{i})dt+\sqrt{\epsilon}dB_{t}.$ (6)
In the gene regulation setting, $c$ could represent CRISPR activation (one element of $c$ being positive) or interference (one element being negative) to enhance or repress gene expression, respectively.
We denote the concatenated intervention column vectors by the matrix $C\in\mathbb{R}^{n\times\mathcal{I}}$. These shifts could correspond to overexpression or knockdown in Perturb-seq data [12]. One appeal of dynamical systems over DAGs is that soft interventions such as these may be expressed very naturally with a shift vector added to the drift in the SDE, while a DAG cannot capture how a feedback loop might be changed through overexpression or knockdown.
In the causal disentanglement literature, soft interventions typically give fewer guarantees [6, 40]. Therefore, we assume knowledge of the interventions,
Assumption 2.2.
The shift interventions $C$ are observed.
3 Related Work
3.1 Causal Representation
Genomics is only one motivation for a formal model of causality [29], a field far too broad to thoroughly survey here. The most popular underlying model is the structural causal model (SCM), which characterizes the conditional distribution of a random variable under arbitrary intervention. Learning an SCM typically requires very strong assumptions such as sparsity [38], interventional data [23], parametric assumptions [30], among many other results. Sparsity is a very common theme in these models, though it may also be expressed in an assumption that the number of latent variables is small, i.e. a low-rank constraint [13].
3.2 Causal Disentanglement with Interventional Data
The bulk of the literature on causal disentanglement focuses on SCMs with an underlying DAG. Particularly relevant to our work are results that assume access to multiple interventional environments, either acting directly on the observed variables [5] or identifying a latent model under some distributional assumptions [23, 40, 6].
Some of these works have addressed the crucial limitation of acyclicity [45, 24, 2], but without necessarily incorporating dynamics. Many works require hard interventions, where an intervened variable is a function of exogenous noise, although some can handle soft interventions [44].
3.3 Dynamical System Methods
Several authors have considered modeling perturbations with dynamical systems [31]. In particular, these models assume observed data is drawn from the stationary distribution under an SDE, and match a learned SDE under numerous interventional regimes. These works focus on linear drift and intervention-dependent parameters [33] and non-linear drift with shift interventions [26]. Notably, neither paper gives theory to confirm whether these interventional models are identifiable.
Nevertheless, works drawing on dynamical systems are regularly applied to the task of inferring GRNs. Some methods act on pseudotime, an inferred notion of time from cell states that enables the machinery of dynamical systems even when very few “real” timepoints are available [42, 18]. Although harder to obtain due the destructive nature of RNA sequencing, genuine temporal data (even with very few timepoints) can also be modeled with the intent of extracting a GRN [25].
The closest work to ours proves an identifiability bound for linear SDEs under a strong sparsity assumption [10]. However, their result doesn’t consider any interventional data, and the exact pattern of sparsity in the drift matrix is assumed to be known a priori, which is rarely the case in applied problems of interest. Another closely related work is [17], which can simultaneously infer the drift and diffusion. However, this work differs from our setting in that it assumes linearity, doesn’t study interventions, and works with multiple temporal marginals rather than just the stationary distribution.
4 Main Results
We consider two main parameterizations, whether the drift is linear or a two-layer neural network (an MLP), and in which regimes the parameters may be uniquely identified up to appropriate invariances. Prior work has considered sparsity in the linear setting only [10], but we focus on parameterizations that project to a low-dimensional space, through small rank or hidden dimension. In other words, although the ambient dimension $n$ may be large, we assume the hidden dimension $r$ is much smaller, and ideally the number of interventions to uniquely identify the parameters should scale with $r$. In the genomics context, $n$ is the number of modeled (e.g., “highly variable”) expressed genes in the thousands, and $r$ corresponds to a much smaller number of co-expressed gene modules/pathways.
4.1 Linear case
We start with the linear case and consider the parameterization $v(x)=(AB-D)x$ where $A\in\mathbb{R}^{n\times r}$, $B\in\mathbb{R}^{r\times n}$, $D\in\mathbb{R}^{n\times n}$. Furthermore, to ensure the matrix is Hurwitz and the SDE therefore has a stationary distribution, we assume $\|AB\|\leq\gamma<1$ and $D\succeq I$. We associate $AB$ with transcriptional regulation, and $D$ with decay of each RNA species.
This parameterization is a simplification of the one used in [33], which also uses linear SDEs but considers hard interventions with intervened genes having entirely new rows in the drift matrix, rather than shift interventions. In our biological context, the low rank assumption is 1) intuitive in that genes are expected to behave in coordinated ways in modules/pathways, and 2) empirically proven, in that almost all successful single-cell methods leverage dimensionality reduction.
It is impossible to get a good deterministic guarantee if the dynamics and interventions are chosen adversarially, as seen in the following proposition:
Proposition 4.1.
In the linear setting, there exist choices for parameters $A,B,D$ and interventions $C$ such that the parameters are not identifiable with less than $n-r$ interventions.
The proof is provided in Appendix A. The intuition is that one can choose interventions that don’t affect the system dynamics. But this situation is pathological, and in the linear case under some weak assumptions about the distribution of the parameters and the interventions we have identifiability. We start with assumptions about the distribution of the SDE parameters and the interventions:
Assumption 4.2.
The matrices $A$ and $B$ are almost surely (a.s.) full-rank, have spectral norm less than one, and are invariant to applying a rotation matrix on the left or the right. Each column of $C$ is drawn iid from some distribution with a density on $\mathbb{R}^{n}$.
The rotational invariance assumption on the distributions of $A$ and $B$ encodes an uninformative prior on which low-rank subspace governs the dynamics. Indeed, a simple choice that satisfies this assumption is sampling $A$ and $B^{T}$ from the uniform measure on the Stiefel manifold (the space of rectangular matrices with orthonormal columns) and then scaling down slightly to ensure a spectral norm strictly smaller than one.
Additionally, for theoretical tractability, we presume knowledge of the decay term $D$. This is a stronger assumption, but defendable in our genomics context. Typically, we’re more interested in dynamics of genes regulating each other rather than self-regulation, and furthermore from a systems perspective there is no way to distinguish self-regulation from RNA decay. While decay rate varies across genes, it is less regulated by other genes than the transcription rate, so external experimental estimates of a diagonal $D$ (e.g., from BRIC-seq [19]) could be used. Hence, fixing the decay rate leaves the focus of learning on the presumed low-rank dynamics.
Assumption 4.3.
The true decay matrix $D$ is observed.
Altogether, we can now present a tight identifiability result in the linear case.
Theorem 4.4.
Consider the linear drift in Equation 6. Suppose $A$, $B$ are drawn from any density further constrained such that $\|A\|,\|B\|\leq\gamma<1$. Then the drift $AB-D$ is identifiable a.s. with $r$ interventions, and unidentifiable a.s. with at most $r-2$ interventions.
See Appendix A for the proof. Naively, one might count $2nr$ unknown parameters in the entries of $A$ and $B$, and assume the $n^{2}$ entries of the stationary covariance $\omega$ would be enough to identify them, but this isn’t the case. One needs exactly enough interventions to account for the hidden rank.
Identifying the decay.
If diagonal decay term $D$ is not inferred in advance, learning it simultaneously is comparable to the setting of robust PCA [8] or recovery of a diagonal plus a positive semidefinite low rank term [37]. However, unlike the usual matrix completion setting where entries of the matrix are revealed uniformly at random, we have access to correlated low-rank measurements of the form $e_{i}c_{j}^{T}$ for $\{e_{i}\}_{i=1}^{n}$ the elementary basis. This setting is well studied with sub-Gaussian measurement vectors [46], but the tools do not readily apply to our setting with non-random measurements and the constraints of the Lyapunov equation. Nevertheless, we observe in Section 5 that identifiability even with an unknown diagonal decay matrix $D$ is empirically achievable in this setting, while still subject to the lower bound established in Theorem 4.4.
4.2 Nonlinear case
The nonlinear case is more challenging, primarily because there is no longer a (nearly) closed form characterization of the stationary distribution. To apply tools from above, we consider a regime where a linearized SDE can approximately capture the true dynamics, by restricting to contractive drift and small noise.
In choosing a sensible drift parameterization, we want to balance expressiveness against low depth, as in single cell applications the induced modules that drive cell evolution are ideally interpretable in their own right, as well as requiring the induced SDE to have a stationary distribution at all.
We consider the vector field $v(x)=A\sigma(Bx)-x$, with the constraints that $\|A\|,\|B\|\leq 1$. Notably, we will not assume prior knowledge of the elementwise map $\sigma:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, only some assumptions:
Assumption 4.5.
The map $\sigma\in\mathcal{C}^{2}$ acts on each element independently and satisfies:
  1. 1.
    $\max_{x\in\mathbb{R}}\sigma_{i}^{\prime}(x)=\gamma<1$
  2. 2.
    $\min_{x\in\mathbb{R}}\sigma_{i}^{\prime}(x)=\tau>0$
  3. 3.
    $\max_{x\in\mathbb{R}}|\sigma_{i}^{\prime\prime}(x)|=M<\infty$
  4. 4.
    The set $\{x\in\mathbb{R}:\sigma_{i}^{\prime\prime}(x)=0\}$ is measure zero.
The strongest constraint here is condition 1, upper bounding the first derivative of the activation, as it implies the noiseless dynamics are globally contractive, but it guarantees a stationary distribution exists for any intervention vector $c$. In practice, cells do not (in fact, cannot) produce unlimited amounts of RNA, so conditions that ensure the existence of a stationary distribution can be motivated from the biological perspective. Furthermore, these conditions still allow for a wide class of possible activations $\sigma$, mainly constrained to be increasing with bounded first and second derivative. We observe the improved expressiveness of the stationary distribution of nonlinear SDEs in Figure 1. Note constraint 4 rules out functions that are linear (or locally linear). This enables a stronger possible identifiability guarantee than only recovering $A$ and $B$ up to rotation as in the linear case.
Figure 1: Contour plot of the stationary SDE under different activations, namely observed linear (top left), intervened linear (bottom left), observed non-linear (top right), and intervened non-linear (bottom right). While the activation contractivity enforces one mode, the linear distribution is Gaussian with fixed covariance across interventions, while the nonlinear distribution can vary more.
Noiseless Setting.
One might be tempted to simply consider the noiseless case, where the SDE reduces to an ordinary differential equation (ODE), and rather than recovering the stationary distribution one recovers instead the unique global stable point. However, the noiseless case cannot take advantage of the low-rank structure, and requires $\Omega(n)$ interventions a.s. for identifiability.
Proposition 4.6.
Setting $\epsilon=0$ in equation 6, the parameters are a.s. not identifiable with fewer than $n-r$ interventions.
The proof is in Appendix A. Intuitively, the equilibrium points of the noiseless ODE approximate the mean of the SDE for small noise (note this is not true for larger noise, for example see [28]). This suggests that in order to make use of the low-rank assumption on the dynamics, one must at least take advantage of second order moments of the SDE even in the small noise limit.
Moments in the zero-noise limit.
As the noise converges to zero, the stationary distribution converges to a dirac centered on the global stable point, and there are no higher order moments. However, the rescaled second-order moments have a non-trivial limit as noise goes to zero:
Theorem 4.7.
Let $x^{*}$ be the unique point satisfying $v(x^{*})=0$, and define $L=Jv(x^{*})$. If $\omega$ is the solution to the Lyapunov equation $L\omega+\omega L^{T}+I=0$, and $m_{\epsilon}$ and $\Sigma_{\epsilon}$ are the mean and covariance of the stationary distribution of Equation (6), we have for sufficiently small $\epsilon$,
$\|m_{\epsilon}-x^{*}\|$ $\leq\left(\frac{\epsilon n}{1-\gamma}\right)^{1/2},$ (7)
$\|\Sigma_{\epsilon}/\epsilon-\omega\|$ $\lesssim\frac{\epsilon^{1/2}r^{2}n^{1/2}\sqrt{M}}{(1-\gamma)^{3}}.$ (8)
The proof is given in Appendix A. This is one instance of perturbation theory for SDEs [15, 35].
Equipped with this fact, one can consider access to the stationary distribution of an SDE, specifically the first and second-order moments, and inspect what happens now as the noise goes to zero. In this case, we obtain non-trivial identifiability, i.e. a bound that scales with the hidden dimension $r$:
Theorem 4.8.
Suppose we observe the first and second moments $m_{\epsilon}$ and $\Sigma_{\epsilon}/\epsilon$ of the stationary distribution of the SDE in Equation (6), in the limit as $\epsilon\rightarrow 0$. Then a.s. with $r+1$ interventions, $A$ and $B^{T}$ are both identifiable up to (simultaneous) permutation and scaling of their columns.
We give a brief sketch of the proof, to capture the novel elements and how the identifiability appears. The zero-noise limit allows us to linearize the SDE under different interventions. However, unlike the truly linear case, the covariance is not identical among all interventions. By Theorem 4.7, the covariance of the $i$th intervention approximately corresponds to the SDE with drift $A(J\sigma(Bx_{i}^{*}))B$ where $x_{i}^{*}$ is the unique zero of the true drift $v(x)+c_{i}$. Sufficient variability in these fixed points enables the identification of $A$ and $B$ by taking linear combinations of these drift matrices. Because the Jacobian term is diagonal, a linear combination of drifts across $r$ interventions can yield a rank-one matrix, which corresponds to simultaneously recovering a column of $A$ and a row of $B$. But we don’t have access to the drift matrix itself, only the covariance it induces, and so the proof relies on finding a different set of matrices where rank-one elements can be identified.
Crucially, one can verify the rank of linear combinations of matrices without knowledge of the activation function itself. This enables the identification of the bias term, but furthermore one can identify the parameters $A$ and $B$ even if each elementwise activation in $\sigma$ is distinct. This observation motivates learnable activations, which we experiment with in Section 5. In our setting where two-layer MLPs with small width offer better interpretability, parameterizing the activation is an effective way to increase expressiveness without sacrificing identifiability.
We hypothesize that the constraint that the SDE drift have a single absorbing point is likely unnecessary. If the ODE given by $\frac{dx}{dt}=v(x)+c_{i}$ had multiple stable equilibria, then in the limit of small noise, the SDE stationary distribution should approach a Gaussian mixture centered around each equilibrium point. The proof only requires certain linear independence conditions of vectors induced by each fixed point, so multiple equilibria and the local covariance around them would offer essentially the same information as additional interventions. This is consistent with the observations in [26] that nonlinear SDE models generalize well with a limited number of interventions.
5 Experiments
To validate the theory given above, we consider experiments demonstrating the recovery of SDE parameters and generalization to unseen interventions. To show the broad applicability of the theory, we consider multiple loss functions: a loss acting directly on the parameters in the linear case [33], a kernelized Stein discrepancy [3], and rollout loss for neural SDEs [21].
5.1 Linear SDE Recovery
Because we are fitting linear SDEs where the stationary distribution has an essentially closed-form (up to solving the Lyapunov equation), we can choose a very simple loss to verify the claim of identifiability (Theorem 4.4). This loss is akin to the one used in Bicycle [33] that fits interventional linear SDEs. Assuming the noise scale $\epsilon$ is known and $L$ denotes the true drift, we fit the parameters $\hat{A}$, $\hat{B}$, and $\hat{D}$. Letting our estimate for the drift be $\hat{L}=\hat{A}\hat{B}-\hat{D}$, the loss function is,
$\mathcal{L}_{lin}(\hat{L})=\|\hat{L}^{-1}C-L^{-1}C\|_{F}+\|\hat{L}\omega+\omega\hat{L}^{T}+\epsilon I\|_{F}.$ (9)
In other words, the loss is an unweighted sum of the mean loss for every intervention, and a loss controlling how well the parameterized drift $\hat{L}$ induces the covariance $\omega$ that is shared among all interventions. We optimize $\hat{A}$, $\hat{B}$ and $\hat{D}$ (constrained to be diagonal) with gradient descent.
We evaluate the recovery of $AB$ under different numbers of interventions for two different ambient dimensions $n$ and two different true ranks $r$ in Figure 2. For each sampled SDE, we use the best train error from 100 independent initializations to deal with the nonconvexity of the objective. The only exception is the oversampled $k=r\log(n)$ where we need only 5 initializations, as the landscape is empirically easier to learn. Further details are given in Section B.
We confirm our theory: when the decay is known, $r-2$ interventions fails to grant identifiability with poor performance and extremely high variance, even with many independent initializations, while $r$ interventions clearly succeed. When the decay isn’t known, we use a modest oversampling of $r\log(n)$ interventions. This amount is informed by the literature on matrix completion, where $\Omega(rn\log(n))$ revealed matrix entries were proven to be necessary for a rank $r$ matrix recovery problem [7]. We scale down by a factor of $n$ because in our setting each intervention yields the mean vector of the perturbed linear SDE in $\mathbb{R}^{n}$, hence $n$ measurements.
(a) Error in $n=100$ with fixed decay
(b) Error in $n=200$ with fixed decay
(c) Error in $n=100$ with learned decay
(d) Error in $n=200$ with learned decay
Figure 2: Normalized Frobenius error of learned drift against true drift in low-rank linear SDEs with $k$ independent Gaussian interventions. Error bars are standard deviation over 5 independent instances.
5.2 Synthetic Nonlinear SDE Generalization
To assess the insight from the theory that the parameters are identifiable with an arbitrary activation function (Theorem 4.8), we propose a simple architectural improvement wherein we learn the activation function. The theory requires the activation function be monotonic, but we consider general learnable MLPs for improved optimization. Furthermore, the simple method for extracting a GRN we apply in section 5.3 requires a two-layer network, so learning the activation is a way to increase expressiveness without sacrificing interpretability. Learnable activations [16] have been proposed before [1] but to our knowledge not previously applied to SDE parameterizations.
To assess generalizability to unseen interventions, we consider a loss function proposed by [26]: the kernel deviation from stationarity (KDS), or equivalently the kernelized Stein discrepancy of the stationary distribution [3]. When the parametric drift is defined as $v_{\theta}(x)=A\sigma(Bx)-Dx$ and the target stationary distribution is $\mu$, the KDS is,
$\mathcal{L}_{KDS}(\theta)=E_{x\sim\mu}E_{x^{\prime}\sim\mu}\mathcal{A}_{x}\mathcal{A}_{x^{\prime}}k(x,x^{\prime}).$ (10)
where $k$ is a given kernel, $\mathcal{A}$ is the generator the parameterized SDE defined as $\mathcal{A}h=v_{\theta}\cdot\nabla h+\frac{\epsilon}{2}\Delta h$, and $\mathcal{A}_{x}$ is the generator acting on the kernel as a function of its first element, likewise $\mathcal{A}_{x^{\prime}}$ acting on the second element. More details on the derivation of this loss are given in [26].
For evaluation, we choose another notion of distributional distance. We observed numerical instability in the Sinkhorn divergence without large entropic regularization [9]. Therefore, we follow [43, 26] and consider the mean squared error (MSE) between the true and predicted distribution. We find training on functions with monotonic activations is numerically unstable when using noise small enough to reach the limit proven in Theorem 4.8, and therefore study more complicated activations in our experiments, using sinusoidal activations. Further details are given in Section B.2.
The results are given in Table 1. We observe a clear benefit for low noise SDEs, which gets smaller as the noise gets bigger and hence further away from our proven settings. This is consistent with our theory, and also intuitive: as the noise gets larger, the stationary distrubtion gets smoother and has less dependence on the intervention vectors. Nevertheless, in the low noise regime we see that learnable activations improve prediction on test interventions without overfitting.
Table 1: Mean distance between true and predicted distribution with $n=20$, $r=3$, averaged over 20 independent seeds with 20 test interventions per seed, comparing a fixed (sigmoid) activation versus a learned activation.
Sigmoid Activation MLP
$\epsilon=0.05$ $21.78\pm 12.78$ $\bm{9.51\pm 1.28}$
$\epsilon=0.1$ $16.54\pm 6.35$ $\bm{9.23\pm 1.65}$
$\epsilon=0.2$ $11.53\pm 3.47$ $\bm{7.96\pm 1.44}$
$\epsilon=0.3$ $7.97\pm 2.19$ $\bm{6.69\pm 2.07}$
5.3 Simulated GRN SDE Generalization
We consider an applied setting for learnable activations, and how they impact the recovery of GRNs. Because there are few fully known regulatory networks we rely on semi-synthetic data to produce data with similar characteristics to true perturbed transcriptomic samples and a ground-truth GRN.
We use the PerturbODE model from [25], similar to our setup and modified to handle a loss for parameterized SDEs [21]. We consider the inferred GRN as the matrix multiplication of the weight matrices in the MLP that parameterizes the drift, corresponding to a first-order Taylor approximation of the dynamics. GRN recovery can be measured as a classification task of individual edges. As a baseline, we also consider the Bicycle interventional linear SDE method [33]. To simulate a known GRN, we use the boolean ODE based-simulator BEELINE [32]. BEELINE uses the provided GRN to parametrize an SDE in the space of mRNA expression and protein expression, with physically plausible regulation dynamics. More precise details are given in the Appendix in Section B.3.
GRN recovery is still extremely difficult (random guessing is $\approx 0.0586$ and maximum AUPRC for dynamical methods $\approx 0.07$), but we observe improvement using nonlinear SDEs over linear SDEs, and substantial improvement using learnable activation in the SDE parameterization (Figure 3).
Figure 3: Gene regulatory network recovery for data from BEELINE among SDE, SDE with learnable activation, and Bicycle model. Error bars are over 5 independent runs.
6 Discussion
Although we can confirm identifiability directly in the linear case because we have access to a simple loss function, direct recovery of parameters is very impractical for the nonlinear SDE parameterization. Nevertheless, the positive results with improved generalization using learned activations is good evidence to confirm the theory. It may be necessary to increase the capacity of the SDE beyond two-layer neural networks, if the extracted GRN could be proven identifiable.
We note some limitations of the the given theory. The results demonstrate identifiability in an idealized setup with infinitesimal small noise and restrictions on the possible drift functions. Generalizing to larger noise regimes may be possible with higher order moment information about the stationary distribution, although in general understanding the stationary distribution outside of linearization is a very challenging task. Considering different parameterizations of the drift or how interventions act on the SDE are especially interesting extensions for future work.
7 Conclusion
In this work, we’ve given the first provable results regarding identifiability of interventional SDEs, in both the linear and nonlinear case. Although such models are currently less popular and less studied then comparable SCM-based modeling, the ability to obtain provable guarantees for dynamical systems without any temporal data suggests they may become more fruitful for causal inference in the future. As longitudinal genomics datasets often have a small number of time-points, future work may consider how to incorporate sparse trajectory information with the stationary distribution for better identifiability guarantees or more effective architectures for recovering regulatory networks.
Acknowledgments
This work was made possible by support from the MacMillan Family and the MacMillan Center for the Study of the Non-Coding Cancer Genome at the New York Genome Center.
Appendix A Proofs of Results
Proof of Proposition 4.1
Consider block matrices such that,
$A$ $=\left[\begin{array}[]{ c | c }\frac{1}{\sqrt{2}}I_{r}&0\\ \hline\cr 0&0\end{array}\right]$ (13)
$B$ $=\left[\begin{array}[]{ c | c }\frac{1}{\sqrt{2}}I_{r}&0\\ \hline\cr 0&0\end{array}\right]$ (16)
and set $D=I$ and $\epsilon=1$. It is then straightforward to confirm these choices yield a Lyapunov equation that satisfies,
$\omega$ $=\left[\begin{array}[]{ c | c }I_{r}&0\\ \hline\cr 0&\frac{1}{2}I_{n-r}\end{array}\right]$ (19)
Furthermore, any skew-symmetric matrix $Q$ obeys the trivial Lyapunov equation $Q\omega+\omega Q^{T}=0$, so we can always add $Q\omega^{-1}$ to the vector field $L$ without changing the covariance.
Choose $Q$ to be only supported in the top left $r\times r$ block, then $AB+Q\omega^{-1}$ will still be rank $r$. And in the worst case, every intervention is of the form $e_{i}$ for $i>r$, and these can’t distinguish between different values of $Q$.
Proof of Theorem 4.4
We show separately the upper and lower bounds for number of interventions needed for almost sure identifiability.
Theorem A.1.
Suppose the true $A$ and $B$ are sampled according to Assumption 4.2, with the constraint that a.s. $\|A\|,\|B\|\leq\sqrt{\gamma}$. Then $L=AB-D$ is a.s. identifiable.
Proof.
Recall that $C=\left[c_{1},\dots,c_{r}\right]$ is the matrix of interventions as columns. The means of the interventional stationary distributions are given by,
$-L^{-1}C=(D^{-1}+D^{-1}A(I-BD^{-1}A)^{-1}BD^{-1})C$ (20)
So with knowledge of $D$ and $C$, we can calculate $A(I-BD^{-1}A)^{-1}BD^{-1}C$. This matrix is a.s. full rank, and its range is equal to the range of $A$. Hence we can infer $P:=P_{A}^{\perp}$.
Now, using the Lyapunov equation, we have,
00 0 $=L\omega P+\omega L^{T}P+\epsilon P$ (21)
$=L\omega P-\omega DP+\epsilon P$ (22)
Rearranging,
$2\omega P$ $=L^{-1}(\omega D+\epsilon I)P$ (23)
$(\omega D+\epsilon I)P$ is rank $n-r$, so a.s. $\operatorname{im}(\omega D+\epsilon I)P\oplus\operatorname{im}C$ is $n$ dimensional. Hence we recover $L^{-1}$ and therefore $L$.
Theorem A.2.
Sample $A\in\mathbb{R}^{n\times r}$ and $B\in\mathbb{R}^{r\times n}$ according to Assumption 4.2. Then a.s. $AB$ is not identifiable with $r-2$ interventions.
Proof.
Assume the decay is fixed at $D=I$, so $L=AB-I$ with induced covariance $\omega$. With $r-2$ interventions, $ABL^{-1}C$ is at most rank $r-2$, and by assumption the kernel of $A^{T}$ is at most dimension $n-r$, so there is guaranteed to be a two dimensional subspace orthogonal to the previous spaces, say with basis vectors $u$ and $v$. Let $Q^{*}=uv^{T}-vu^{T}$.
Then we consider $\hat{L}=L+\omega Q$ where $Q=B^{T}A^{T}Q^{*}AB$, and claim that $\hat{L}$ is a valid, distinct drift matrix that generates the same data.
First, note that $u,v$ are orthogonal to the kernel of $A^{T}$ and $B^{T}$ is full-rank a.s., so $B^{T}A^{T}Q^{*}$ is non-zero. Transposing and applying the same reasoning, we get that $Q=B^{T}A^{T}Q^{*}AB\neq 0$. Scaling down the magnitude of $Q^{*}$ if necessary, we have that $\hat{L}$ is Hurwitz and distinct from $L$.
Second, note that $\hat{L}=(I+\omega B^{T}A^{T}Q^{*})AB-I$, so it satisfies the rank $r$ constraint on the non-diagonal part.
Now we show it agrees on the induced stationary distributions. Again from the choice of the two dimensional subspace, we have that $Q^{*}ABL^{-1}C=0$, which implies $QL^{-1}C=0$.
From the Woodbury matrix identity, we see the means of all stationary distributions are conserved,
$\hat{L}^{-1}C$ $=(L+\omega Q)^{-1}C$ (24)
$=\left[L^{-1}-L^{-1}\omega(I+QL^{-1}\omega)QL^{-1}\right]C$ (25)
$=L^{-1}C$ (26)
Finally, note that by antisymmetry of $Q$,
$\hat{L}\omega+\omega\hat{L}^{T}$ $=L\omega+\omega L^{T}+\omega Q\omega+\omega Q^{T}\omega$ (27)
$=L\omega+\omega L^{T}$ (28)
which implies $\hat{L}$ satisfies the same Lyapunov equation and therefore induces the same covariance $\omega$. ∎
Proof of Proposition 4.6
In the zero-noise setting, the SDE reduces to an ODE, and in the limit as time goes to infinity, the stationary distribution is simply a point mass on the unique fixed point of the drift. Therefore, one only gets access to the stationary points $x_{i}^{*}$ for each vector field $v+c_{i}$: other higher moments are not defined.
Suppose we have $n-r-1$ interventions. Since $A$ has $r$ columns, there must be a vector $u$ in $\mathbb{R}^{n}$ orthogonal to the columns of $A$ and $C$.
Let $\hat{B}=B+bu^{T}$ for any non-trivial $b\in\mathbb{R}^{r}$. We claim $\hat{B}$ induces the same data as $B$. First, note that because $u$ is orthogonal to every $c_{i}$, and $x_{i}^{*}-c_{i}$ is in the image of $A$, it follows $u$ is also orthogonal to each $x_{i}^{*}$.
Thus, one can confirm that $\hat{B}x_{i}^{*}=Bx_{i}^{*}$, so the fixed points don’t change. But $\hat{B}\neq B$, and they are clearly not equivalent up to permutation or scaling invariance.
Proof of Theorem 4.7
We will require two lemmas.
Lemma A.3.
Consider a vector field $v(x)=A\sigma(Bx)-x+c$ with $\|A\|,\|B\|\leq 1$, $\|\sigma\|_{\infty}=\gamma<1$. Let $p$ be the stationary distribution of our usual SDE, $x^{*}$ be the unique stationary point of $v$, and $P$ be an orthogonal projection such that $BP=B$ and $PA=A$. Assume $2(k-1)\leq Tr(P)$, then,
$E_{p}\left[\|P(x-x^{*})\|^{2k}\right]$ $\leq\left(\frac{\epsilon Tr(P)}{1-\gamma}\right)^{k}.$ (29)
Proof.
Let $f(x)=A\sigma(Bx)+c$. Note $f$ is a contraction and $f(x)=f(Px)$. We have by Cauchy-Schwarz,
$P(x-x^{*})\cdot v(x)$ $=P(x-x^{*})\cdot(v(x)-v(x^{*}))$ (30)
$=P(x-x^{*})\cdot(f(x)-f(x^{*})-(x-x^{*}))$ (31)
$=P(x-x^{*})\cdot(f(Px)-f(Px^{*})-(x-x^{*}))$ (32)
$\leq-(1-\gamma)\|P(x-x^{*})\|^{2}$ (33)
Choose $g(x)=\|P(x-x^{*})\|^{2k}$, then if $T:=Tr(P)$ we have,
$\nabla g_{k}(x)$ $=2kg_{k-1}(x)P(x-x^{*})$ (34)
$\Delta g_{k}(x)$ $=2kTg_{k-1}(x)+4k(k-1)g_{k-1}(x)$ (35)
The generator of the SDE is $\mathcal{A}g=\nabla g\cdot v+\frac{\epsilon}{2}\Delta g$. So Fokker-Planck gives,
00 0 $=E_{p}\left[\mathcal{A}g_{k}\right]$ (36)
$=E_{p}\left[\nabla g_{k}\cdot w(x)+\frac{\epsilon}{2}\Delta g_{k}\right]$ (37)
$=E_{p}\left[2kg_{k-1}P(x-x^{*})\cdot v(x)+\frac{\epsilon}{2}(2kT+4k(k-1))g_{k-1}(x)\right]$ (38)
$\leq E_{p}\left[-2k(1-\gamma)g_{k}+\frac{\epsilon}{2}(2kT+4k(k-1))g_{k-1}(x)\right]$ (39)
Simple algebra gives,
$E_{p}\left[g_{k}\right]$ $\leq\frac{\epsilon(T+2(k-1))}{2(1-\gamma)}E_{p}\left[g_{k-1}(x)\right]$ (41)
Now apply the assumption $2(k-1)\leq T$ and the result follows by induction. ∎
Lemma A.4.
Assume same conditions as previous lemma. Then the Taylor expansion with remainder around $x^{*}$ given by
$v(x)=Jv(x^{*})(x-x^{*})+R(x),$ (42)
satisfies the bound
$\|R(x)\|\leq 2r\sqrt{2M}\|P(x-x^{*})\|^{2},$ (43)
where $M:=\sup_{i}\|\sigma_{i}^{\prime\prime}\|_{\infty}$.
Proof.
Assume that $\operatorname{im}A\oplus(\ker B)^{\perp}$ is only supported on the first $2r$ coordinates. The $i$th coordinate of the remainder must take the form,
$R_{i}(x)$ $=v_{i}(x)-e_{i}^{T}Jv(x^{*})(x-x^{*})$ (44)
$=\sum_{|\alpha|=2}\frac{2}{\alpha!}(x-x^{*})^{\alpha}\int_{0}^{1}(1-t)\partial_{\alpha}v_{i}(x^{*}+t(x-x^{*}))dt$ (45)
By assumption, second order derivatives of $v$ are always bounded by $M$, and they are identically zero if $i>2r$. This implies $R_{i}$ is only nonzero for $i\leq 2r$, in which case,
$|R_{i}(x)|$ $\leq M\sum_{|\alpha|=2,\alpha_{>2r}=0}\frac{2}{\alpha!}(x-x^{*})^{\alpha}$ (46)
$\leq 4rM\|P(x-x^{*})\|^{2}$ (47)
where $P$ the projection onto the first $2r$ coordinates.
Hence,
$\|R(x)\|^{2}\leq 8r^{2}M\|P(x-x^{*})\|^{4}.$ (48)
Now, drop the assumption that $\operatorname{im}A\oplus(\ker B)^{\perp}$ is restricted to the first $2r$ elements.
Choosing an appropriate orthogonal matrix $Q$, it’s easy to see that $\tilde{v}(x):=Q^{T}v(Qx)$ satisfies the above assumption, with a fixed point $y^{*}=Q^{T}x^{*}$, and that if we let $y=Qx$, then
$\|v(x)-Jv(x^{*})(x-x^{*})\|=\|\tilde{v}(y)-J\tilde{v}(\tilde{y}^{*})(y-\tilde{y}^{*})\|$ (49)
so the bound follows for $v$ as well.
Using these results, we can proceed to the main perturbation theory result:
Proof of Theorem 4.7.
We will drop the $\epsilon$ subscript, where it is clear that $m:=m_{\epsilon}$ and $\Sigma:=\Sigma_{\epsilon}$ are the first and second order moments of the stationary distribution.
The mean bound follows from Lemma A.3 and Jensen’s inequality,
$\|m-x^{*}\|$ $=\|E_{p}[x-x^{*}]\|$ (50)
$\leq E\left[\left\|x-x^{*}\right\|^{2}\right]^{1/2}$ (51)
$\leq\left(\frac{\epsilon n}{1-\gamma}\right)^{1/2}.$ (52)
Fokker-Planck yields a formula for second-order moments of linear SDE stationary distribution (see for example Chapter 5.5 in [36]), such that,
00 0 $=E_{p}[(x-m)v(x)^{T}+v(x)(x-m)^{T}+\epsilon I]$ (53)
Note the simple fact that
$E_{p}[(x-m)(x-x^{*})^{T}]$ $=E_{p}[(x-m)(x-m+m-x^{*})^{T}]$ (54)
$=E_{p}[(x-m)(x-m)^{T}]=\Sigma$ (55)
combined with the linearization of $v$ we can write,
00 0 $=E_{p}[(x-m)(x-x^{*})^{T}L^{T}+L(x-x^{*})(x-m)^{T}+\epsilon I+(x-m)R(x)^{T}+R(x)(x-m)^{T}]$ (56)
$=L\Sigma+\Sigma L^{T}+\epsilon I+E_{p}[(x-m)R(x)^{T}+R(x)(x-m)^{T}]$ (57)
and dividing by $\epsilon$,
$0=L\frac{\Sigma}{\epsilon}+\frac{\Sigma}{\epsilon}L^{T}+I+\frac{1}{\epsilon}E_{p}[(x-m)R(x)^{T}+R(x)(x-m)^{T}]$ (58)
As in the lemmas, we let $P$ be the orthogonal projection onto $\operatorname{im}A\oplus(\ker B)^{\perp}$. Then by Lemma A.3, Lemma A.4 and Cauchy-Schwarz we have,
$E_{p}\left[\|x-m\|\cdot\|R(x)\|\right]$ $\lesssim r\sqrt{M}E_{p}\left[(\|x-m\|\cdot\|P(x-x^{*})\|^{2}\right]$ (59)
$\lesssim r\sqrt{M}E_{p}\left[\|x-m\|^{2}\right]^{1/2}E_{p}\left[\|P(x-x^{*})\|^{4}\right]^{1/2}$ (60)
$\lesssim\frac{\epsilon r^{2}\sqrt{M}}{1-\gamma}E_{p}\left[\|x-m\|^{2}\right]^{1/2}$ (61)
$\lesssim\frac{\epsilon r^{2}\sqrt{M}}{1-\gamma}E_{p}\left[\|x-x^{*}\|^{2}+\|x^{*}-m\|^{2}\right]^{1/2}$ (62)
$\lesssim\frac{\epsilon^{3/2}r^{2}n^{1/2}\sqrt{M}}{(1-\gamma)^{2}}$ (63)
By the fact that $\|xy^{T}\|\leq\|x\|\cdot\|y\|$ and Jensen’s inequality,
$\left\|\frac{1}{\epsilon}E_{p}\left[(x-m)R(x)^{T}+R(x)(x-m)^{T}\right]\right\|$ $\lesssim\frac{1}{\epsilon}E_{p}\left[\|x-m\|\cdot\|R(x)\|\right]$ (64)
$\lesssim\frac{\epsilon^{1/2}r^{2}n^{1/2}\sqrt{M}}{(1-\gamma)^{2}}$ (65)
Now, if we choose $\epsilon$ small enough to guarantee the above bound is strictly less than 1, then the Lyapunov equation given in Equation (58) has a unique solution. It’s easy to then check from the integral form of the solution that,
[…]
By assumption, $L=A\sigma^{\prime}B-I$ is Hurwitz, and in fact $\|L\|<1-\gamma$. By the fact that $\|e^{L}\|\leq e^{\|L\|}$,
[…]
$\lesssim\frac{\epsilon^{1/2}r^{2}n^{1/2}\sqrt{M}}{(1-\gamma)^{3}}$ (69)
Proof of Theorem 4.8
We first need a result about the equilibrium points of the drift, ruling out any degeneracies. As before, we let $x_{i}^{*}$ denote the unique zero of $v(x)+c_{i}=A\sigma(Bx)-x+c_{i}$.
Lemma A.5.
For $r$ independently sampled interventions, $x_{i}^{*}-c_{i}$ are a.s. linearly independent for $1\leq i\leq r$.
Proof.
Let $g(y,z)=\sigma(BAy+z)$. By the contractive constraint on the activation, and the constraints that $\|A\|,\|B\|\leq 1$, it’s clear $g$ is a uniform contraction mapping. The derivatives of $g$ satisfy,
$J_{y}g(y,z)$ $=\sigma^{\prime}(BAy+z)BA$ (70)
$J_{z}g(y,z)$ $=\sigma^{\prime}(BAy+z)$ (71)
with all spectral norms bounded by $\gamma<1$.
By the uniform contraction mapping principle, the fixed point map $y^{*}(z)$ such that $g(y^{*}(z),z)=y^{*}(z)$ is differentiable, with Jacobian
$Jy^{*}(z)=(I-J_{y}g(y^{*}(z),z))^{-1}J_{z}g(y^{*}(z),z)$ (72)
Notice this Jacobian never vanishes, and furthermore $\|y^{*}(z)\|\rightarrow\infty$ as $\|z\|\rightarrow\infty$. This follows from the fact that $\sigma^{\prime}\geq\tau$,
$\|y^{*}(z)\|$ $=\|\sigma(BAy^{*}(z)+z)\|$ (73)
$\geq\|\sigma(0)+\tau(BAy^{*}(z)+z)\|$ (74)
which is clearly impossible for $\|z\|$ sufficiently large and $\|y^{*}\|$ bounded.
Therefore, by the Hadamard global inverse theorem, $y^{*}(z)$ is a diffeomorphism on $\mathbb{R}^{r}$.
Now, consider the random variable $y^{*}(Bc)$ where $c$ is a sampled intervention vector. Conditioning on $B$ being full-rank, $Bc$ has a density in $\mathbb{R}^{r}$, and therefore so does $y^{*}(Bc)$.
The terms $y^{*}(Bc_{i})$ are sampled independently and hence are a.s. linearly independent. By uniqueness of fixed points, it must be the case that $x_{i}^{*}-c_{i}=Ay^{*}(Bc_{i})$. Therefore, given $r$ interventions, the vectors $x_{i}^{*}-c_{i}$ for $1\leq i\leq r$ are also a.s. linearly independent.
Lemma A.6.
The set of vectors $\sigma^{\prime}(Bx_{0}^{*})/(\sigma^{\prime}(Bx_{i}^{*}))-\mathbf{1}$, where /// denotes elementwise division, for $1\leq i\leq r$ are a.s. a basis for $\mathbb{R}^{r}$.
Proof.
By the same argument as in Lemma A.5, we have that $\sigma(Bx_{i}^{*})=y^{*}(Bc_{i})$, and therefore since $\sigma$ is smooth and strictly monotonic, we have that each $Bx_{i}^{*}=\sigma^{-1}(y^{*}(Bc_{i}))$ has a density and is independent of the other interventions.
Now, conditioning on the value of $\sigma^{\prime}(Bx_{0}^{*})$, we want to show that $\sigma^{\prime}(Bx_{0}^{*})/(\sigma^{\prime}(Bx_{i}^{*}))-\mathbf{1}$ for each index $i$ are linearly independent.
We’ll proceed by induction. Suppose the first set of $i$ such vectors are linearly independent, and consider the $i+1$th vector.
Let $H$ be the hyperplane formed by the first $i$ vectors. Define $f(z)=\sigma^{\prime}(Bx_{0}^{*})/(\sigma^{\prime}(z))-\mathbf{1}$, then we equivalently want to show that $P(f(Bx_{i+1}^{*})\in H)=0$. It’s quick to see that $Jf(z)$ is diagonal for any $z$, and rank deficient only when there’s some index $j$ such that $\sigma_{j}^{\prime\prime}(z)$ is zero. By assumption on $\sigma$, this set of points is measure zero. Hence the Jacobian of $f$ is a.s. full-rank. And it is folklore that $f$ is therefore necessarily a nonsingular map, i.e., its preimage maps measure zero sets to measure zero sets (for a proof see [34]). Finally, because $H$ is measure zero, we deduce $P(Bx_{i+1}^{*}\in f^{-1}(H))=0$ since $Bx_{k+1}^{*}$ has a density.
Lemma A.7.
Suppose $X\in\mathbb{R}^{n\times r}$ has full rank and $Y\in\mathbb{R}^{r\times r}$ is invertible. Furthermore, suppose the matrices $D_{1},\dots,D_{r}\in\mathbb{R}^{r\times r}$ are diagonal, with the vectors along each diagonal linearly independent.
Then only observing $XD_{i}Y$ for $1\leq i\leq r$, $X$ and $Y$ are uniquely identifiable up to permutation and scaling of their columns and rows, respectively.
Proof.
First, note that by the independence assumption, there’s some choice of $\alpha\in\mathbb{R}^{r}$ such that $\sum_{i=1}^{r}\alpha_{i}D_{i}$ has only one non-zero element along its diagonal, and therefore $\sum_{i=1}^{r}\alpha_{i}XD_{i}Y$ is rank-one.
Conversely, suppose $\sum_{i=1}^{r}\alpha_{i}XD_{i}Y$ were rank-one but $\sum_{i=1}^{r}\alpha_{i}D_{i}$ had at least two non-zero elements on the diagonal, w.l.o.g. the first and second elements. We will index $X$ by its columns and $Y$ by its rows. By the independence assumptions, there are vectors $u_{1},u_{2}$ such that $y_{i}\perp u_{1}$ iff $i\neq 1$, and likewise for $u_{2}$. Thus,
$\left(\sum_{i=1}^{r}\alpha_{i}XD_{i}Y\right)u_{1}$ $=X\left(\sum_{i}\alpha_{i}D_{i}\right)e_{1}\langle y_{1},u_{1}\rangle=\beta_{1}x_{1}$ (75)
$\left(\sum_{i=1}^{r}\alpha_{i}XD_{i}Y\right)u_{2}$ $=X\left(\sum_{i}\alpha_{i}D_{i}\right)e_{2}\langle y_{2},u_{2}\rangle=\beta_{2}x_{2}$ (76)
where by assumption $\beta_{1},\beta_{2}\neq 0$. Again by independence, $x_{1}$ and $x_{2}$ are not collinear contradicting the rank-one assumption.
Thus, a linear combination of terms that is rank-one offers a certificate that the given linear combination $\alpha$ does in fact yield $\sum_{i=1}^{r}\alpha_{i}D_{i}$ with only one non-zero term. By considering all possible $\alpha$ vectors, one will arrive at each possible rank-one term $\gamma_{i}x_{i}y_{i}^{T}$ for some unknown scalar $\gamma_{i}$. ∎
For the proof of the main theorem, because we have a total of $r+1$ interventions we will zero-index the interventions and consider $0\leq i\leq r$ for convenience. We introduce the notation $\sigma_{[i]}^{\prime}$ to denote the Jacobian of $\sigma$ evaluated at $Bx_{i}^{*}$, to distinguish it from an element of $\sigma$. Because $\sigma$ acts elementwise, $\sigma_{[i]}^{\prime}$ is necessarily a diagonal matrix.
Proof of Theorem 4.8.
By Theorem 4.7, for each intervention in the limit of zero noise we recover the mean vector $x_{i}^{*}$ and the covariance matrix $\omega_{i}$, which satisfies $Jv(x_{i}^{*})\omega_{i}+\omega_{i}(Jv(x_{i}^{*}))^{T}+I=0$. The Jacobian of the vector field satisfies $Jv(x_{i}^{*})=A\sigma_{[i]}^{\prime}B-I$.
From Lemma A.5, any $r$ terms $x_{i}^{*}-c_{i}$ are linearly independent. Since they are all in the image of $A$, we can derive $P_{A}$ and hence $P_{A}^{\perp}$. From a similar calculation as in the linear case, one can confirm that,
$2\omega_{i}P_{A}^{\perp}=(I-\frac{1}{2}A\sigma_{[i]}^{\prime}B)^{-1}P_{A}^{\perp},$ (77)
and taking the transpose,
$2P_{A}^{\perp}\omega_{i}=P_{A}^{\perp}(I-\frac{1}{2}B^{T}\sigma_{[i]}^{\prime}A^{T})^{-1}.$ (78)
This matrix is rank $n-r$, so its kernel is dimension $r$ and clearly generated by the columns of
$(I-\frac{1}{2}B^{T}\sigma_{[i]}^{\prime}A^{T})A$ (79)
Thus, taking an arbitrary basis of the above mentioned kernel must be of the form,
$Z_{i}:=(I-\frac{1}{2}B^{T}\sigma_{[i]}^{\prime}A^{T})AQ_{i},$ (80)
for some invertible matrix $Q_{i}\in\mathbb{R}^{r\times r}$. Note that here, $Z_{i}$ is observed but $Q_{i}$ is not.
Note that,
$P_{A}^{\perp}Z_{i}=-\frac{1}{2}P_{A}^{\perp}B^{T}\sigma_{[i]}^{\prime}A^{T}AQ_{i}.$ (81)
Observe two facts: 1) $\sigma_{[i]}^{\prime}A^{T}AQ_{i}$ is a.s. an invertible matrix in $\mathbb{R}^{r\times r}$, and 2) $P_{A}^{\perp}B^{T}$ maps from $\mathbb{R}^{r}$ to an $n-r$ subspace, so if $n>2r$ then a.s. this is full-rank.
Because $P_{A}^{\perp}Z_{i}$ and $P_{A}^{\perp}Z_{0}$ have the same range, we can derive the unique $M_{i}$ such that,
$P_{A}^{\perp}Z_{i}M_{i}=P_{A}^{\perp}Z_{0}.$ (82)
Furthermore the pseudoinverse $(P_{A}^{\perp}B^{T})^{\dagger}$ acts as a left inverse, so we derive the identity,
$\sigma_{[i]}^{\prime}A^{T}AQ_{i}M_{i}=\sigma_{[0]}^{\prime}A^{T}AQ_{0}.$ (83)
Notice that $Z_{i}$ and $M_{i}$ are observed variables while the $Q_{i}$ are not. Nevertheless, we can rewrite,
$Z_{i}M_{i}-Z_{0}$ $=AQ_{i}M_{i}-AQ_{0}$ (84)
$=A(A^{T}A)^{-1}\left(A^{T}AQ_{i}M_{i}-A^{T}AQ_{0}\right)$ (85)
$=(A^{\dagger})^{T}\left(\sigma_{[0]}^{\prime}(\sigma_{[i]}^{\prime})^{-1}-I\right)A^{T}AQ_{0}$ (86)
By Lemma A.6, the diagonal parts of $\left(\sigma_{[0]}^{\prime}(\sigma_{[i]}^{\prime})^{-1}-I\right)$ are linearly independent for $1\leq i\leq r$.
Now, by Lemma A.7, we observe the rows of $(A^{\dagger})^{T}$ and the columns of $A^{T}AQ_{0}$ up to permutation and scaling. W.l.o.g. we can assume the permuted rows of $(A^{\dagger})^{T}$ are placed in the true order, and we can conclude we observe $(A^{\dagger})^{T}\Lambda_{1}^{-1}$ and $\Lambda_{2}A^{T}AQ_{0}$ for some unobserved, diagonal scaling matrices $\Lambda_{1}$ and $\Lambda_{2}$.
Taking transpose and another pseudoinverse, we recover $A\Lambda_{1}$, so it remains to recover $B$.
We can observe,
$(-2P_{A}^{\perp}Z_{i})(\Lambda_{2}A^{T}AQ_{0})^{-1}$ $=P_{A}^{\perp}B^{T}\sigma_{[i]}^{\prime}A^{T}AQ(\Lambda_{2}A^{T}AQ_{0})^{-1}$ (87)
$=P_{A}^{\perp}B^{T}\sigma_{[i]}^{\prime}\Lambda_{2}^{-1}.$ (88)
Taking a transpose, we’ve observed $\Lambda_{2}^{-1}\sigma_{[i]}^{\prime}BP_{A}^{\perp}$.Now, returning to equation 77, we can rewrite the observed matrix with Woodbury,
$2\omega_{i}P_{A}^{\perp}$ $=(I-\frac{1}{2}A\sigma_{[i]}^{\prime}B)^{-1}P_{A}^{\perp}$ (89)
$=\left(I+\frac{1}{2}A\left(I-\frac{1}{2}\sigma_{[i]}^{\prime}BA\right)^{-1}\sigma_{[i]}^{\prime}B\right)P_{A}^{\perp}$ (90)
Subtracting $P_{A}^{\perp}$, applying the left inverse of $A\Lambda_{1}$ and the right inverse of $\Lambda_{2}^{-1}\sigma_{[i]}^{\prime}B^{T}P_{A}^{\perp}$ and finally multiplying by two, we observe,
$\Lambda_{1}^{-1}(I-\frac{1}{2}\sigma_{[i]}^{\prime}BA)^{-1}\Lambda_{2}$ (91)
The whole matrix is $\mathbb{R}^{r\times r}$ and invertible, we so observe,
$\Lambda_{2}^{-1}\Lambda_{1}-\frac{1}{2}\Lambda_{2}^{-1}\sigma_{[i]}^{\prime}BA\Lambda_{1}$ (92)
Subtracting from a different index $j$ and rescaling, we get,
$\Lambda_{2}^{-1}(\sigma_{[i]}^{\prime}-\sigma_{[j]}^{\prime})BA\Lambda_{1}$ (93)
Notice that
$((A\Lambda_{1})^{T}A\Lambda_{1})^{-1}(A\Lambda_{1})^{T}=\Lambda_{1}^{-1}(A^{T}A)^{-1}A^{T},$ (94)
and so multiplying this term on the right, we observe
$\Lambda_{2}^{-1}(\sigma_{[i]}^{\prime}-\sigma_{[j]}^{\prime})BP_{A}.$ (95)
Because we observe $\Lambda_{2}^{-1}\sigma_{[i]}^{\prime}BP_{A}^{\perp}$ as above, we can subtract across two indicies and get $\Lambda_{2}^{-1}(\sigma_{[i]}^{\prime}-\sigma_{[j]}^{\prime})BP_{A}^{\perp}$. Finally, because $P_{A}+P_{A}^{\perp}=I$, we at last observe $\Lambda_{2}^{-1}(\sigma_{[i]}^{\prime}-\sigma_{[j]}^{\prime})B$. Hence we’ve also recovered $B$ up to scaling of its rows (as long as $(\sigma_{[i]}^{\prime}-\sigma_{[j]}^{\prime})$ has no zeros along the diagonal, which is true a.s.).
Appendix B Experimental Details
B.1 Linear SDE Recovery
For the true simulated linear SDE, $A$ and $B$ are sampled with iid Gaussian entries, and then normalized to have spectral norm equal to $0.9$. The true decay matrix $D$ has each diagonal entry sampled iid from the uniform distribution on the interval $[1,2]$. The model is trained with the loss given in Equation (9), with the Adam optimizer [22] with an initial learning rate of $0.005$ and $3000$ iterations. Additionally, due to the non-convexity of the objective, on each training instance we run 100 independent initializations and pick the one with the smallest training error. This excludes the oversampled setting with $k=r*\log(n)$, where we get good performance with only 5 independent initializations.
The plotted mean and standard deviations in Figure 2 are based on 5 separate instantiations of the true model and fitting a new model on the stationary distribution. Experiments were run on CPU. All experiments (including those below) were run on a Linux system.
B.2 Synthetic Nonlinear SDE Generalization
[…]
We consider a fixed true SDE, with hidden dimension $r=3$, where the rectangular matrices $A$ and $B^{T}$ have one down the diagonal and zero elsewhere, $D=I$, $\gamma=0.98$, and the elementwise activation function is given by
$\sigma_{1}(x)$ $=3\cos(3(x-0.5))$
$\sigma_{2}(x)$ $=2\sin(2(x+1.5))-1$
$\sigma_{3}(x)$ $=\sigma_{1}(x)$
The hidden dimension of the activation MLPs is fixed at 20.
Each intervention is sampled as a Gaussian vector with mean zero and variance $0.1$ on each entry, and from each intervened SDE we collect 5000 samples. To do this, we use the Euler-Maruyama scheme to solve the SDE, and use MCMC to draw approximately independent samples from the stationary distribution, using $dt=0.01$, a thinning factor of $300$ and the first 500 samples treated as burn-in. We use an radial basis function kernel to parameterize the KDS loss.
We train with 10 interventions, and evaluate on 20 held-out interventions. Training is done with AdamW [27], with initial learning rate $0.003$ and 50000 iterations. The hyperparameter ranges considered are given in Table 2. All experiments were run on CPU.
Table 2: Hyperparameter ranges considered for synthetic nonlinear SDE experiments.
Hyperparameter Range
model hidden size $r$$\{4,8,16\}$
kernel bandwidth $\{3,5,7\}$
$L_{1}$ weight regularization $\{0,10^{-5},10^{-4}\}$
B.3 Simulated GRN SDE Generalization
We first give a summary of the simulator. BoolODE is scRNA-seq data simulator that samples mRNA readouts via a stochastic differential equation (SDE). The underlying cyclic GRN is encoded in the SDE via a Hill function based approximation to a boolean logical circuit. $x_{i}$, $p_{i}$, and $R[i]$ each represents the mRNA concentration for gene $i$, the concentration of the protein encoded by gene $i$, and the set of proteins that regulate gene $i$. Further, $r$, $l_{p}$, $l_{x}$, and $s$ are scalar coefficients representing translation rate, RNA decay rate, protein decay rate, and noise standard deviation respectively. The functions $f_{i}(\cdot)$ encodes the gene regulatory network. The default SDE is given by,
$\frac{dx_{i}}{dt}$ $=f_{i}(p_{R[i]})-l_{x}x_{i}+s\sqrt{x_{i}}dB_{t}$ (96)
$\frac{dp_{i}}{dt}$ $=rx_{i}-l_{p}p_{i}+s\sqrt{p_{i}}dB^{\prime}_{t}$ (97)
where $B$ and $B^{\prime}$ are independent Brownian motions.
We simulate overexpression (e.g., CRISPR-a) experiments by inducing perfect intervention coupled with increased transcription for a set of intervened genes $I_{k}$. For each intervened gene $j\in I_{k}$ we set $f_{j}(\cdot)=0$ and add a positive shift (set to 20 in the experiments) to $dx_{j}/dt$. We consider $K$ such intervention regimes, plus the observational setting ($k=0$) with no intervention, i.e, $I_{0}=\emptyset$, for a total of $K+1$ regimes. With this setup, we obtain a family of distributions $\{\rho_{k}\}_{k=0}^{K}$ in gene expression space.
We adapt the neural ODE architecture proposed in [25]. The base SDE model under $I_{k}$ parametrizes the change of gene expression via the drift function as below. $A$ and $B$ are the coefficient matrices encoding the modular graph. $\alpha$ is a scaling vector controlling the rate of non-linear activation in the module, where the activation $\sigma$ is the logistic sigmoid. Further, $\beta$ is a bias term that shifts the activation threshold of the modules. The GRN is extracted as $A\,\text{diag}(\alpha\circ\mathbf{1}_{N})B$. The intervention is modeled as a combination of standard basis vectors, $\sum_{j\in I_{k}}e_{j}$, specifying the overexpression. The model also learns a single diffusion coefficient scalar. Lastly, $M$ is a masking matrix blocking the signals from the intervened genes’ regulators. Altogether, the drift is given by,
$v_{k}(x)=M_{k}A\sigma(\alpha\circ(Bx-\beta))+\sum_{j\in I_{k}}e_{j}-Dx.$ (98)
We alternatively consider the architecture using a learnable MLP to model the nonlinear activation of modular signals,
$v_{k}(x)=M_{I_{k}}A\sigma_{*}(Bx-\beta)+\sum_{j\in I_{k}}e_{j}-Dx$ (99)
where $\sigma_{*}$ denotes the MLP as described above. The model loss is the Sinkhorn divergence [14] applied to the true perturbed distribution and the samples drawn from the learned SDE.
Each interventional distribution $\rho_{k}$ is obtained by taking the observational data $\rho_{0}$ as initial distribution and simulating the SDE via the Euler-Maruyama method.
We fit the hyperparameters by testing the ODE architecture in [25] on the same simulated datasets, and then doing zero-shot transfer of the hyperparameters to the SDE models. For Bicycle, we sweep the hyperparameters in Table 3 and use the package defaults for the remaining values. Experiments are run on an nvidia tesla V100 gpu.
Table 3: Hyperparameter ranges for GRN simulation.
Hyperparameter Range
scale $L_{1}$$\{0.0001,0.001,0.01,0.1,1.0\}$
scale spectral $\{0,1.0\}$
scale Lyapunov $\{0.1,1,10\}$
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The theory results claimed by the abstract are stated in the body and proved in section A, with the experimental details given in section B.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations of the method and its application are given in the discussion in section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The full proof of all results and subsequent assumptions are given in section A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Architectures and hyperparameter sweeps are given for each experiment in section B.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The supplement contains all code used to produce the results and figures in the paper body.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Hyperparameter sweeps, optimizers and other such choices are given in section B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All experiments are accompanied either by error bars or by standard deviation over a specified number of independent runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Resources used to run the experiments are mentioned in section B
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read and followed the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The work is mainly theory based; although long-term study of gene regulatory inference has social implications, we believe this work is an early effort in this application of machine learning and doesn’t require a more elaborate analysis of social impact at this time.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All libraries and datasets used are cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: There are no new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: There are no crowdsourced experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: There are no study participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were not used.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.